# Exact Random Graph Matching with Multiple Graphs

### v1

@@ Abstract
This work studies fundamental limits for recovering the underlying correspondence among multiple correlated random graphs. We identify a necessary condition for any algorithm to correctly match all nodes across all graphs, and propose two algorithms for which the same condition is also sufficient. The first algorithm employs global information to simultaneously match all the graphs, whereas the second algorithm first partially matches the graphs pairwise and then combines the partial matchings by transitivity. Both algorithms work down to the information theoretic threshold. Our analysis reveals a scenario where exact matching between two graphs alone is impossible, but leveraging more than two graphs allows exact matching among all the graphs. Along the way, we derive independent results about the $k$-core of Erdős-Rényi graphs.
@@ 1 Introduction
The information age has ushered an abundance of correlated networked data. For instance, the network structure of two social networks such as Facebook and Twitter is correlated because users are likely to connect with the same individuals in both networks. This wealth of correlated data presents both opportunities and challenges. On one hand, information from various datasets can be combined to increase the fidelity of data - translating to better performance in downstream learning tasks. On the other hand, the interconnected nature of this data also raises privacy and security concerns. Linkage attacks, for instance, exploit correlated data to identify individuals in an anonymized network by linking to other sources [NS09]. This poses a significant threat to user privacy.
Graph matching is the problem of recovering the underlying latent correspondence between correlated networks. The problem finds many applications in machine learning: de-anonymizing social networks [NS08, NS09], identifying similar functional components between species by matching their protein-protein interaction networks [BSI06, KHGPM16], object detection [SS05] and tracking [YYL+16] in computer vision, and textual inference for natural language processing [HNM05]. In most applications of interest, data is available in the form of several correlated networks. For instance, social media users are active each month on 6.7 social platforms on average [Ind23]. Similarly, reconciling protein-protein interaction networks among multiple species is an important problem in computational biology [SXB08]. As a first step toward this objective, many research works have studied the problem of matching two correlated graphs.
@@ 1.1 Related Work
The theoretical study of graph matching algorithms and their performance guarantees has primarily focused on Erdős-Rényi (ER) graphs. Pedarsani and Grossglauser [PG11] introduced the subsampling model to generate two such correlated graphs. The model entails twice subsampling each edge independently from a parent ER graph to obtain two sibling graphs, both of which are marginally ER graphs themselves. The goal is then to match nodes between the two graphs to recover the underlying latent correspondence. This has been the framework of choice for many works that study graph matching. For example, Cullina and Kiyavash studied the problem of exactly matching two ER graphs, where the objective is to match all vertices correctly [CK16, CK17]. They identified a threshold phenomenon for this task: exact recovery is possible if the problem parameters are above a threshold, and impossible otherwise. Subsequently, threshold phenomena were also identified for partial graph matching between ER graphs - where the objective is to match only a positive fraction of nodes [GML21, HM23, WXY22, DD23]. The case of almost-exact recovery - where the objective is to match all but a negligible fraction of nodes - was studied by Cullina and co-authors: a necessary condition for almost exact recovery was identified, and it was shown that the same condition is also sufficient for the $k$-core estimator [CKMP19]; the estimator is described formally in Section 3. This estimator proved useful to uncover the fundamental limits for graph matching in other contexts such as the stochastic block model [GRS22] and inhomogeneous random graphs [RS23]. Ameen and Hajek [AH23] showed some robustness properties of the $k$-core estimator in the context of matching ER graphs under node corruption. The estimator plays an important role in the present work as well.
A sound understanding of ER graphs inspires algorithms for real-world networks. Various efficient algorithms have been proposed, including algorithms based on the spectrum of the graph adjacency matrices [FMWX22], node degree and neighborhood based algorithms [DCKG19, DMWX21, MRT23] as well as algorithms based on iterative methods [DL23] and counting subgraphs [MWXY23, BCL+19]. Some of these are discussed in Section 5 in relation to the present work.
Incorporating information from multiple graphs to match them has been recognized as an important research direction, for instance in the work of Gaudio and co-authors [GRS22]. To our knowledge, the only other work to consider matchings among multiple graphs is the work of Rácz and Sridhar [RS21]. However, the authors solve a different problem of matching stochastic block models, and note that it is possible to exactly match $m$ graphs whenever it is possible to exactly match any two graphs by pairwise matching all the graphs exactly. In contrast, we show that under appropriate conditions, it is possible to exactly match $m$ ER graphs even when no two graphs can be pairwise matched exactly.
Contributions
In this work, we investigate the problem of combining information from multiple correlated networks to boost the number of nodes that are correctly matched among them. We consider the natural generalization of the subsampling model to generate $m$ correlated random graphs, and identify a threshold such that it is impossible for any algorithm to match all nodes correctly across all graphs when the problem parameters are below this threshold. Conversely, we show that exact recovery is possible above the threshold. This characterization generalizes known results for exact graph matching when $m=2$. Subsequently, we show that there is a region in parameter space for which exactly matching any two graphs is impossible using only the two graphs, and yet exact graph matching is possible among $m>2$ graphs using all the graphs.
We present two algorithms and prove their optimality for this task. The first algorithm matches all $m$ graphs simultaneously based on global information about the graphs. In contrast, the second algorithm first pairwise matches graphs, and then combines them to match all nodes across all graphs. We show that both algorithms correctly match all the graphs all the way down to the information theoretic threshold. Finally, we illustrate through simulation that our subroutine to combine information from pairwise comparisons between networks works well when paired with efficient algorithms for graph matching. Our analysis also yields some theoretical results about the $k$-core of ER graphs that are of independent interest.
2 Preliminaries and Setup
@@
In this work, $G\sim\mathsf{ER}(n,p)$ denotes that the graph $G$ is sampled from the Erdős-Rényi distribution with parameters $n$ and $p$, i.e. $G$ has $n$ nodes and each edge is independently present with probability $p$. For a graph $G$, we denote the set of its vertices by $V\equiv V(G)$ and its edges by $E(G)$. The edge status of each vertex pair $\{i,j\}$ with $i\neq j$ is denoted by $G\{i.j\}$, so that $G\{i,j\}=1$ if $\{i,j\}\in E(G)$ and $G\{i,j\}=0$ otherwise. The degree of a node $v$ in graph $G$ is denoted $\delta_{G}(v)$. Let $\pi$ denote a permutation on $V(G)=\{1,\cdots,n\}$. For a graph $G$, denote by $G^{\pi}$ the graph obtained by permuting the nodes of $G$ according to $\pi$, so that
$G\{i,j\}=G^{\pi}\left\{\pi(i),\pi(j)\right\}\ \forall\ i,j\in V(G)\text{ such that }i\neq j.$
Standard asymptotic notation $(O(\cdot),o(\cdot),\cdots)$ is used throughout and it is implicit that $n\to\infty$.
Subsampling model
Consider the subsampling model for correlated random graphs [PG11], which has a natural generalization to the setting of $m$ graphs. In this model, a parent graph $G$ is sampled from the Erdős-Rényi distribution $\mathsf{ER}(n,p)$. The $m$ graphs $G_{1},G_{2}^{\prime},\cdots,G_{m-1}^{\prime},G_{m}^{\prime}$ are obtained by independently subsampling each edge from $G$ with probability $s$. Finally, the graphs $G_{2},\cdots,G_{m}$ are obtained by permuting the nodes of each of the graphs $G_{2}^{\prime},\cdots,G_{m}^{\prime}$ respectively according to independent permutations $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ sampled uniformly at random from the set of all permutations on $[n]$, i.e.
$G_{j}=(G_{j}^{\prime})^{\pi^{*}_{1j}}\ \text{ for all }\ j\in\{2,\cdots,m\}.$
Figure 1 illustrates this process of obtaining correlated graphs using the subsampling model. In this work, we are interested in the setting where $s$ is constant and $p=C\log(n)/n$ for some $C>0$.
Objective 1.
Determine conditions on parameters $C$, $s$ and $m$ so that given correlated graphs $G_{1},\cdots,G_{m}$ from the subsampling model, it is possible to exactly recover the underlying correspondences $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ with probability $1-o(1)$.
Stated thus, the underlying correspondences use the graph $G_{1}$ as a reference. Thus, for ease of notation, we will use $G_{1}$ and $G_{1}^{\prime}$ interchangeably. Note that the underlying correspondence between all the graphs is fixed upon fixing $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$: for any two graphs $G_{i}$ and $G_{j}$, their underlying correspondence is given by $\pi^{*}_{ij}:=\pi^{*}_{1j}\circ(\pi^{*}_{1i})^{-1}$.
Formally, a matching $(\mu_{12},\cdots,\mu_{1m})$ is a collection of injective functions with domain $\mathsf{dom}(\mu_{1i})\subseteq V$ for each $i$, and co-domain $V$. An estimator is simply a mechanism to map any collection of graphs $(G_{1},\cdots,G_{m})$ to a matching. We say that an estimator completely matches the graphs if the output mappings $\mu_{12},\cdots\mu_{1m}$ are all complete, i.e. they are all permutations on $\{1,\cdots,n\}$.
Figure 1: Illustration of obtaining $m$ correlated graphs from the subsampling model
3 Main Results and Algorithm
This section presents necessary and sufficient conditions to meet 1.
Theorem 2 (Impossibility).
Let $G_{1},\cdots,G_{m}$ be correlated graphs obtained from the subsampling model with parameters $C$ and $s$, and let $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ denote the underlying latent correspondences between $G_{1}$ and $G_{2},\cdots,G_{m}$ respectively. Suppose that
$Cs\left(1-(1-s)^{m-1}\right)<1.$
The output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ of any estimator satisfies
$\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\ \widehat{\pi}_{13}=\pi^{*}_{13},\cdots,\ \widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=o(1).$
Theorem 2 implies that the condition $Cs(1-(1-s)^{m-1}>1$ is a necessary condition to exactly match $m$ graphs with probability bounded away from 00. We show that this condition is also sufficient to exactly match $m$ graphs with probability going to $1$.
Theorem 3 (Achievability).
Let $G_{1},\cdots,G_{m}$ be correlated graphs obtained from the subsampling model with parameters $C$ and $s$, and let $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ denote the underlying latent correspondences between $G_{1}$ and $G_{2},\cdots,G_{m}$ respectively. Suppose that
$Cs\left(1-(1-s)^{m-1}\right)>1.$

### v2

# Aligning Multiple Inhomogeneous Random Graphs: Fundamental Limits of Exact Recovery

@@ Abstract
This work studies fundamental limits for recovering the underlying correspondence among multiple correlated graphs. In the setting of inhomogeneous random graphs, we present and analyze a matching algorithm: first partially match the graphs pairwise and then combine the partial matchings by transitivity. Our analysis yields a sufficient condition on the problem parameters to exactly match all nodes across all the graphs. In the setting of homogeneous (Erdős-Rényi) graphs, we show that this condition is also necessary, i.e. the algorithm works down to the information theoretic threshold. This reveals a scenario where exact matching between two graphs alone is impossible, but leveraging more than two graphs allows exact matching among all the graphs. Converse results are also given in the inhomogeneous setting and transitivity again plays a role. Along the way, we derive independent results about the $k$-core of inhomogeneous random graphs.
@@ 1 Introduction
The information age has ushered an abundance of correlated networked data. For instance, the network structure of two social networks such as Facebook and Twitter is correlated because users are likely to connect with the same individuals in both networks. This wealth of correlated data presents both opportunities and challenges. On one hand, information from various datasets can be combined to increase the fidelity of data, which translates to better performance in downstream tasks. On the other hand, the interconnected nature of this data also raises privacy and security concerns. Linkage attacks, for instance, exploit correlated data to identify individuals in an anonymized network by linking to other sources (NS, 09).
The problem of graph alignment is to recover the underlying latent node correspondence between correlated networks, i.e. to match the nodes across graphs. The problem derives its importance from the fact that nodes in networks are often unlabeled or scrambled – for example, user identities in social networks are anonymized for privacy. Applications include:
  1. 1.
    Social networks: Matching two social networks amounts to recovering the hidden node identities of users in the anonymized network. Narayanan and Shmatikov (NS, 09) established the efficacy of this approach: A third of the users active on both Twitter and Flickr could be identified with only a 12% error rate, simply by matching the anonymous Twitter graph to the Flickr graph. In a related work (NS, 08), they also showed that anonymous movie ratings of Netflix users can be recovered by matching to the public Internet Movie Database (IMDb).
  2. 2.
    Biological networks: The protein-protein interactomes of two different species are often correlated networks. Singh and co-authors (SXB, 08) showed that matching the PPI network of yeast to that of fruit flies allowed identification of conserved functional components between the two species. This transfer of knowledge between species is a useful tool in biology and genomics (KHGPM, 16; BSI, 06). Similarly, the brain connectomes of healthy humans are correlated (STK, 05), and their alignment is useful in detecting abnormalities (CPPDG, 24).
  3. 3.
    Natural language processing: Given a passage of text and a hypothesis, consider the objective of determining whether the hypothesis is entailed by the passage. Haghighi and co-authors (HNM, 05) presented an approach to solving this, by first constructing dependency trees of the text and hypothesis respectively, and then matching them to conclude with a yes/no answer after thresholding a score of the output matching.
  4. 4.
    Computer vision: Graph matching is employed extensively for vision tasks. One example is object detection: Schellewald and Schnörr (SS, 05) showed that objects in images can be detected by matching graph representations of the test image to a feature image that contains the object. Yet another example in pattern recognition is the problem of tracking the movement of objects in a sequence of frames (YYL+, 16).
  5. 5.
    Autonomous agents: Robots and self-driving cars rely on a multitude of sensors to collect data about their environment. Combining information from various sensors increases data fidelity. For example, Ye and co-authors (YDF+, 22) showed that graph alignment is useful for scene-map construction when multiple robots explore an unknown environment.
There is a long history of utilizing domain knowledge to implement graph alignment algorithms for these tasks – see for example (YYL+, 16; VF, 13). Even so, the theoretical study of graph matching algorithms and their performance guarantees is a relatively recent undertaking. The vast majority of work has focused on pairs of random graphs, particularly Erdős-Rényi graphs. Presently, we study exact recovery thresholds with multiple inhomogeneous random graphs.
@@ 1.1 Related Work
The theoretical study of graph matching algorithms and their performance guarantees has focused on the homogeneous setting of Erdős-Rényi (ER) graphs. Pedarsani and Grossglauser (PG, 11) introduced the subsampling model to generate two such correlated graphs: the model entails twice subsampling each edge independently from a parent ER graph to obtain two sibling graphs, both of which are marginally ER graphs themselves. The goal is then to match nodes between the two graphs to recover the underlying latent correspondence. This has been the framework of choice for many works that study graph matching. For example, Cullina and Kiyavash (CK, 16, 17) studied the problem of exactly matching two ER graphs, where the objective is to match all vertices correctly. They identified a threshold phenomenon for this task: exact recovery is possible if the problem parameters are above a threshold, and impossible otherwise. Subsequently, threshold phenomena were also identified for partial graph matching between ER graphs, where the objective is to match only a positive fraction of nodes (GML, 21; HM, 23; WXY, 22; DD, 23; Du, 25). The case of almost-exact recovery – where the objective is to match all but a negligible fraction of nodes – was studied by Cullina and co-authors (CKMP, 19): a necessary condition for almost exact recovery was identified, and the same condition was shown to also sufficient be for the $k$-core estimator; the estimator is described in Section 3. This estimator proved useful for graph matching in other contexts such as the stochastic block model (GRS, 22) and inhomogeneous random graphs (RS, 23). Ameen and Hajek (AH, 23) showed some robustness properties of the $k$-core estimator in the context of matching ER graphs under node corruption. The estimator also plays an important role in the present work.
Other works have studied computational aspects of graph alignment: Various efficient algorithms have been proposed, including algorithms based on the spectrum of the graph adjacency matrices (FMWX, 22), node degree and neighborhood based algorithms (DCKG, 19; DMWX, 21; MRT, 23) as well as algorithms based on iterative methods (DL, 23) and counting subgraphs (MWXY, 23; BCL+, 19; CR, 24). While all these works study Erdős-Rényi graphs, there is recent interest in computational aspects of graph alignment in other models such as the stochastic block model (CR, 24), as well as inhomogeneous random graphs (DFW, 23). Still other works have studied the graph matching problem under side information – this includes settings where a subset of nodes are matched a priori and provided as seeds (PG, 11; MX, 20; LS, 18), the setting where node attributes are present (WZWW, 24; ZWWW, 24; YC, 24), and the setting where correlation is localized to a subset of nodes (HSY, 24).
Incorporating information from multiple graphs to match them has been recognized as an important research direction, for instance in the work of Gaudio and co-authors (GRS, 22). For example, a user is active, on average, on 6.7 social networks each month (Ind, 23). Similarly, reconciling protein-protein interaction networks among multiple species is an important problem in computational biology (SXB, 08). These applications also necessitate going beyond ER graphs, since many real-world networks lack the homogeneity of the ER model. To our knowledge, a very limited number of papers consider matchings among multiple graphs: The works (JLK, 21) and (RS, 21) have different objectives, and note that it is possible to exactly match $m$ graphs whenever it is possible to exactly match any two graphs by pairwise matching all the graphs exactly. In contrast, we show that under appropriate conditions, it is possible to exactly match $m$ graphs even when no two graphs can be pairwise matched exactly. The work of Rácz and Zhang (RZ, 24) studies the interplay between graph matching and community detection for multiple correlated stochastic block models. The recent work of Vassaux and Massoulié (VM, 25) studies the information-theoretic limit for exact recovery in the Gaussian model, where the observations are multiple correlated Wigner matrices, and also the limit for partial recovery in the homogeneous Erdős-Rényi setting. Closest to our work is the work of Rácz and Sridhar (RS, 23), where sufficeint conditions for exact and partial graph matching between two inhomoegenous random graphs are studied.
1.2 Contributions
The present work studies fundamental limits of exact recovery with multiple inhomogeneous random graphs. These graphs allow for heterogeneity by having each edge $\{i,j\}$ in the graph appear independently with probability $p_{ij}$, and subsume other models such as ER graphs and stochastic block models, and are closely related to random geometric graphs.
We consider the generalization of the matching problem to $m$ graphs and propose an algorithm to combine information among the graphs. A key idea is that of transitive closure – our algorithm works by pairwise matching graphs partially and then boosting the matchings by transitivity. Our analysis yields a sufficient condition for exact recovery of multiple inhomogeneous graphs. In the homogeneous case where $p_{ij}=p$, we show that this condition is also necessary, thereby characterizing the sharp reconstruction threshold for exactly matching multiple Erdős-Rényi graphs. In the heterogeneous setting, we also provide converse results and transitive closure plays a role here as well. We present simulation results to corroborate the utility of transitive closure as a black-box, efficient and optimal bridge between the binary and $m$-ary matching problems.
2 Model
Let $\mathbf{P}=(p_{ij})_{i,j\in[n]}$ with $0\leq p_{ij}=p_{ji}\leq 1$ denote a symmetric matrix with zero diagonals. A graph $G$ is an inhomogeneous random graph on the vertex set $[n]\triangleq\{1,2,\cdots,n\}$ with parameter matrix $\mathbf{P}$, denoted by $G\sim\mathrm{RG}(n,\mathbf{P})$, if each edge $\{i,j\}$ with $i<j$ exists independently with probability $p_{ij}$. Consider the natural extension of the subsampling model to $m$ graphs. Let $s\in(0,1]$, and obtain $m$ correlated inhomogeneous graphs $G_{1}^{\prime},\cdots,G_{m}^{\prime}$ by subsampling each edge of a parent graph $G\sim\mathrm{RG}(n,\mathbf{P})$ independently with probability $s$. These sibling graphs are themselves inhomogeneous random graphs with parameter matrix $\mathbf{P}s$. Let $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ denote independent permutations that are each drawn uniformly at random from the set of all permutations on $[n]$. Obtain the graphs $G_{2},\cdots,G_{m}$ by permuting the nodes of $G_{k}^{\prime}$ according to $\pi^{*}_{1k}$, for each $k=2,\cdots,m$. For example, $\{i,j\}$ is an edge in $G_{k}^{\prime}$ if and only if $\{\pi^{*}_{1k}(i),\pi^{*}_{1k}(j)\}$ is an edge in $G_{k}$. An illustration of the subsampling model is provided in Figure 1.
Figure 1: Obtaining $m$ correlated inhomogeneous graphs via the subsampling model.
We will use the tuple $(\pi_{12},\cdots,\pi_{1m})$ of permutations to identify the correspondence between all $m$ graphs. Stated thus, the graph $G_{1}$ is used as a reference and so we will interchangeably use $G_{1}$ and $G_{1}^{\prime}$. Note that the underlying correspondence between all the graphs is fixed upon fixing $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$: for any two graphs $G_{k}$ and $G_{\ell}$, their correspondence is given by $\pi^{*}_{k\ell}\triangleq\pi^{*}_{1\ell}\circ(\pi^{*}_{1k})^{-1}$.
In this work, we are concerned with exact recovery. The goal is to determine necessary and sufficient conditions on the parameters $\mathbf{P}$, $s$ and $m$ so that given correlated inhomogeneous graphs $G_{1},\cdots,G_{m}$ from the subsampling model, it is possible to exactly recover the underlying correspondences $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ with probability $1-o(1)$.
Formally, a matching $(\mu_{12},\cdots,\mu_{1m})$ is a collection of injective functions with domain $\mathrm{dom}(\mu_{1i})\subseteq V$ for each $i$, and co-domain $V$. An estimator is simply a mechanism to map any collection of graphs $(G_{1},\cdots,G_{m})$ to a matching. We say that an estimator completely matches the graphs if the output mappings $\mu_{12},\cdots\mu_{1m}$ are all complete, i.e. they are all permutations on $[n]$. If the matching $(\mu_{12},\cdots,\mu_{1m})$ satisfies $\mu_{1j}=\pi^{*}_{1j}$ for all $2\leq j\leq m$, then all nodes are correctly matched across all graphs, i.e. the graphs are exactly matched.
Some special cases of inhomogeneous random graphs
The inhomogeneous random graph model captures a variety of generative models. Following (RS, 23), we introduce these models conditioned on latent variables intrinsic to them for convenience.
  1. 1.
    Erdős-Rényi graph: Given a parameter $p\in[0,1]$, the homogeneous case of ER graphs sets $p_{ij}=p$ for all distinct $i$ and $j$.
  2. 2.
    Stochastic block model (SBM) (Abb, 18; HLL, 83): Given a partition $V_{1},\cdots,V_{M}$ of the node set $[n]$ into $M$ communities and edge probabilities $(q_{ab})_{1\leq a<b\leq M}$, the SBM sets $p_{ij}=q_{ab}$ for all distinct $i$ and $j$ such that $i\in V_{a}$ and $j\in V_{b}$.
  3. 3.
    Random geometric graph (RGG) (LR, 23): Given a collection $x_{1},\cdots,x_{n}$ of points in a set $S\subseteq\mathbb{R}^{d}$, and parameters $p\in[0,1]$ and $r>0$, the RGG sets $p_{ij}=p\cdot\mathbf{1}_{\left\{\|x_{i}-x_{j}\|_{2}\leq r\right\}}$ for distinct $i$ and $j$.
  4. 4.
    Chung-Lu graph (CLG) (CL, 02): Given a collection $(w_{i})_{1\leq i\leq n}$ of positive weights satisfying $w_{i}\leq\sqrt{\sum_{j\in[n]}w_{j}}$ for all $i$, the CLG model sets $p_{ij}=\frac{w_{i}w_{j}}{\sum_{k\in[n]}w_{k}}$ for distinct $i$ and $j$.
Correlated graphs from any of the above models are obtained via the subsampling model.
@@ 2 Model
Throughout this work, $p_{\max}$ denotes the largest entry in the matrix $\mathbf{P}$, i.e. $p_{\max}=\max_{i,j}p_{ij}$. For a graph $G$ and node $v$ in $G$, let $\delta_{G}(v)$ denote the degree of $v$. Given a matching $\widehat{\mu}$, we denote by $G^{\mu}$ the graph on node set given by the image of $\mu$ obtained by relabeling the nodes of $G$ according to $\mu$, i.e. for $i,j\in\mathrm{dom}(\mu)$:
$\{i,j\}\in E(G)\iff\{\mu(i),\mu(j)\}\in E(G^{\mu}).$
For any two graphs $H_{1}$ and $H_{2}$ on the same vertex set $V$, denote by $H_{1}\vee H_{2}$ their union graph and by $H_{1}\wedge H_{2}$ their intersection graph. An edge $\{i,j\}$ is present in $H_{1}\vee H_{2}$ if it is present in either $H_{1}$ or $H_{2}$. Similarly, the edge is present in $H_{1}\wedge H_{2}$ if it is present in both $H_{1}$ and $H_{2}$. For a random variable $X$, denote by $M_{X}(t)\triangleq\mathbb{E}[e^{tX}]$ the moment generating function of $X$. Standard asymptotic notation $(O(\cdot),o(\cdot),\cdots)$ is used throughout, and it is implicit that $n\to\infty$.
3 Main Results and Algorithms
3.1 Exact recovery threshold: Positive results
Our first result is a sufficient condition for exactly matching all nodes across all the graphs.
Theorem 1
Let $d_{i}=\sum_{j\neq i}p_{ij}$. Suppose that there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p_{\max}=o(n^{\alpha/2-\varepsilon-1})$ and
$\sum_{i=1}^{n}e^{-d_{i}s(1-(1-s)^{m-1})}=o(n^{-\alpha}).$ (1)